IN A SAMPLE OF OWLS IT IS FOUND THAT . There were 35 red males, 70 red females, 50 gray males, and 45 gray females IS THE COLOR INDEPENDANT OF SEX OF SAMPLE



1
Expert's answer
2022-02-21T16:50:18-0500

The data is incomplete.

Question:

Find the probability if a red male is selected.

Solution:

There were 35 red males, 70 red females, 50 gray males, and 45 gray females.

Total = 35+70+50+45=200

Required probability = "\dfrac{35}{200}=\dfrac7{40}"